OpenStudy (anonymous):

Write the equation of a line in slope intercept form that is perpendicular to the line y = –4x and passes through the point (2, 6)

OpenStudy (anonymous):

@Conqueror @thomas5267 @Spyro517 @mybabygirl34 @iGreen @jigglypuff314 @Joel_the_boss

OpenStudy (bloomlocke367):

do you know how to write an equation that's perpendicular?

OpenStudy (bloomlocke367):

You flip the slope and change the sign... but that's only part of the answer. Let me know what you get and I'll work you through the rest

OpenStudy (bloomlocke367):

@HWBUSTER00

OpenStudy (anonymous):

is it y-6=(-4)(x-2) @BloomLocke367

OpenStudy (bloomlocke367):

not quite...

OpenStudy (bloomlocke367):

you didn't flip the slope. instead of -4, it should be -1/4

OpenStudy (bloomlocke367):

oops 1/4 should be positive

OpenStudy (anonymous):

oh so y-6=(1/4)(x-2) @BloomLocke367

OpenStudy (bloomlocke367):

Yes, I'm pretty sure that's correct :) Hang on, I'll double check it for you

OpenStudy (bloomlocke367):

Yep :) Do you need to put it in the form y=mx+b? or can you leave it like that?

OpenStudy (bloomlocke367):

@HWBUSTER00

OpenStudy (anonymous):

nope just answered the question. But thanks!!!!! @BloomLocke367

OpenStudy (bloomlocke367):

No problem :)